Company Z-prime’s earnings and dividends per share are expected to grow by 4% a year. Its growth will stop after year 4. In year 5 and afterward, it will pay out all earnings as dividends. Assume next year’s dividend is $2, the market capitalization rate is 12% and next year’s EPS is $9. What is Z-prime’s stock price?
Answer #1

We need to follow the below steps to calculate the price:

  1. Calculate the earnings of year 5
  2. Calculate terminal value
  3. Calculate the PV of all dividends and terminal value
  4. Sum all the PV to find the price of the stock
Particulars 1 2 3 4 5
Earnings $       9.00 $       9.36 $       9.73 $    10.12 $    10.53
Dividends $       2.00 $       2.08 $       2.16 $       2.25 $    10.53

So year 5 earnings are 10.53 at 4% growth rate each year.

The entire earnings from year 5 onwards is distributed as dividends so it is a perpetuity as the growth rate is 0

PV of perpetuity at the end of year 4 or beginning of year 5 = Dividend/capitalization rate

PV of perpetuity at the end of year 4 or beginning of year 5 = 10.53/0.12 = $87.74

Now we find the PV of the dividends and terminal value:

Year CF Discount Factor Discounted CF
1 $      2.00 1/(1+0.12)^1= 0.892857143 0.892857142857143*2= $      1.79
2 $      2.08 1/(1+0.12)^2= 0.797193878 0.79719387755102*2.08= $      1.66
3 $      2.16 1/(1+0.12)^3= 0.711780248 0.711780247813411*2.1632= $      1.54
4 $      2.25 1/(1+0.12)^4= 0.635518078 0.635518078404831*2.249728= $      1.43
4 $   87.74 1/(1+0.12)^4= 0.635518078 0.635518078404831*87.739392= $    55.76
Price = Sum of all Discounted CF $    62.17

So the price of the stock = $62.17